A Complete Bibliography of the Journal of Statistical Physics: 1960–1969

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

31 May 2024 Version 1.03

Title word cross-reference

acoustic [Sei69]. answers [Bal69, Cla69]. applications [Sch69]. approach [HR69, KtH69].

between [Maz69a]. body [Maz69c]. Boltzmann [HD69]. Book [Fut69]. boundary [Cer69]. Brownian [Maz69a, Maz69b, Maz69c].

change [BC69]. collision [NDR69]. computer [SZ69]. continuum [BN69]. contours [NDR69]. correctness [Tru69]. coupled [HR69]. covariance [And69]. critical [NH69].

dependence [Gro69]. derivation [BC69]. detection [Hel69]. Differential [PR69]. dilute [OB69]. disks [OB69, SS69]. displacement [NH69]. distribution [Maz69c]. DNA [SZ69]. DNA-like [SZ69]. dynamic [Mor69]. effect [ELK69]. electrodynamics [GK69]. electromagnetic [Sei69]. energy [Gro69]. energy-entropy [Gro69]. engineering [Sch69]. Entropy [HR69, OB69, Gro69, PR69]. equation [BC69, HD69]. equations [BC69]. equilibrium [HR69]. estimation [Hel69]. exciton [ELK69].

ferromagnet [KtH69]. fluctuations [BN69]. function [KtH69, Maz69c]. functions [GK69].

gas [HD69, OB69]. Gaussian [Mee69]. General [BC69]. generation [And69]. geometry [SS69]. Green [KtH69].

hard [OB69]. harmonic [HR69]. Heisenberg [KtH69]. helix [SZ69]. hydrodynamic [BC69].

II [Maz69b]. III [Maz69c]. incoherent [Sei69]. information [Hob69, HR69, Rei69d]. integrals [NDR69, Sch69]. Interaction [Maz69a]. Introduction [Rei69a]. inverse [And69]. ionized [BC69]. irreversibility [OB69].

kinetic [Cer69, Tru69].

lattice [ELK69]. learning [Gro69]. lifetimes [ELK69]. like [Rei69d, SZ69]. limit [Tru69]. limited [ELK69]. Limiting [SS69]. linear [GK69].

Mathematical [Mor69]. Maxwell [BC69]. mechanics [GK69, Ric69]. modeling [Mor69]. molecules [BC69]. motion [Maz69a, Maz69b, Maz69c].

Navier [Tru69]. nearly [Mee69]. nearly-Gaussian [Mee69]. networks [Gro69]. News [Ano69a, Ano69b, Ano69c]. nonlinear [Mee69, Mor69]. nonrecurrent [Gro69]. Nonuniform [Maz69b]. Note [NH69]. nucleation [Kik69].

oscillator [HR69].

paradox [Kik69]. paramagnetic [KtH69]. particles [Maz69a]. Pattern [Ric69]. phase [KtH69]. physics [Sch69]. plasma [GK69]. point [NH69]. polyatomic [HD69]. polytope [SS69]. precise [Tru69]. principle [Cer69]. problem [And69]. problems [Cer69, Mor69]. processes [Mee69]. Prospectus [Rei69b, Rei69c].

Quantum [Hel69, NH69]. Questions [Bal69, Cla69].

recognition [Ric69]. recurrent [Gro69]. representation [Mee69]. respect [Tru69]. response [GK69]. review [Fut69]. rigid [SS69]. rods [SS69]. rotation [Kik69].

scattering [Sei69]. signed [Gro69]. simulation [SZ69]. space [NDR69]. spheres [SS69]. stationary [And69]. statistical [BC69, GK69, Ric69]. Stochastic [Mee69, Mor69]. Stokes [Tru69]. superconductivity [Gin69]. Superfluidity [Gin69]. system [BC69]. systems [BN69, HR69, Maz69b, Mor69].

theorem [Hob69]. theory [BN69, Cer69, Hel69, Hob69, HR69, Kik69, Maz69a, Maz69b, Maz69c, Rei69d, Tru69]. Thermodynamic [Rei69d]. Thermodynamic-like [Rei69d]. tiling [PR69]. transformations [Rei69d]. translation [Kik69]. translation-rotation [Kik69]. trap [ELK69]. trap-limited [ELK69]. Two [Maz69c]. Two-body [Maz69c].

universe [Gin69]. unwinding [SZ69]. upper [Tru69].

value [Cer69]. variational [Cer69]. Velocity [NDR69]. Velocity-space [NDR69]. vibrations [ELK69].

wave [Sei69]. Wiener [Sch69].

References

Anderson:1969:IPS

[And69] Brian D. O. Anderson. The inverse problem of stationary covariance generation. Journal of Statistical Physics, 1(1):133–147, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

         URL http://link.springer.com/article/10.1007/BF01007246.

Anonymous:1969:Na

[Ano69a] Anonymous. News. Journal of Statistical Physics, 1(1):229–230, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

         (electronic). URL http://link.springer.com/article/10.1007/
         BF01007252.

Anonymous:1969:Nb

[Ano69b] Anonymous. News. Journal of Statistical Physics, 1(3):519, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

         (electronic). URL http://link.springer.com/article/10.1007/
         BF01106584.

Anonymous:1969:Nc

[Ano69c] Anonymous. News. Journal of Statistical Physics, 1(4):597–599, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.1007/
BF01024135.

Balakrishnan:1969:QA

[Bal69] A. V. Balakrishnan. Questions and answers. Journal of Statistical Physics, 1(1):227–228, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01007251.

Beshinske:1969:SDH

[BC69] R. J. Beshinske and C. F. Curtiss. A statistical derivation of the hydrodynamic equations of change for a system of ionized molecules. I. General equation of change and the Maxwell equations. Journal of Statistical Physics, 1(1):163–174, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01007248.

Babloyantz:1969:TFC

[BN69] Agnessa Babloyantz and G. Nicolis. On the theory of fluctuations in continuum systems. Journal of Statistical Physics, 1(4):563–570, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.1007/
BF01024131.

Cercignani:1969:VPB

[Cer69] Carlo Cercignani. A variational principle for boundary value problems in kinetic theory. Journal of Statistical Physics, 1(2):297–311, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.1007/
BF01007482.

Clark:1969:QA

[Cla69] Michael G. Clark. Questions and answers. Journal of Statistical Physics, 1(4):595, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.com/

article/10.1007/BF01024134.

Elliott:1969:ELV

[ELK69] Richard A. Elliott, Katja Lakatos, and Robert S. Knox. The effect of lattice vibrations on trap-limited exciton lifetimes. Journal of Statistical Physics, 1(2):253–269, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01007480.

Futrelle:1969:BR

[Fut69] R. P. Futrelle. Book review. Journal of Statistical Physics, 1(3):517, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.1007/
BF01106583.

Ginzburg:1969:SSU

[Gin69] V. L. Ginzburg. Superfluidity and superconductivity in the universe. Journal of Statistical Physics, 1(1):3–24, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:

//link.springer.com/article/10.1007/BF01007238.

Golden:1969:ESM

[GK69] K. I. Golden and G. Kalman. The electrodynamics and statistical mechanics of linear plasma response functions. Journal of Statistical Physics, 1(3):415–466, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.com/

article/10.1007/BF01106580.

Grossberg:1969:LEE

[Gro69] Stephen Grossberg. On learning and energy-entropy dependence in recurrent and nonrecurrent signed networks. Journal of Statistical Physics, 1(2):319–350, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01007484.

Hoffman:1969:BEP

[HD69] David K. Hoffman and John S. Dahler. The Boltzmann equation for a polyatomic gas. Journal of Statistical Physics, 1(4):521–558, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.1007/
BF01024129.

Helstrom:1969:QDE

[Hel69] Carl W. Helstrom. Quantum detection and estimation theory. Journal of Statistical Physics, 1(2):231–252, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01007479.

Hobson:1969:NTI

[Hob69] Arthur Hobson. A new theorem of information theory. Journal of Statistical Physics, 1(3):383–391, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01106578.

Huerta:1969:EIT

[HR69] Manuel A. Huerta and Harry S. Robertson. Entropy, information theory, and the approach to equilibrium of coupled harmonic oscillator systems. Journal of Statistical Physics, 1(3):393–414, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/BF01106579.

Kikuchi:1969:TRP

[Kik69] Ryoichi Kikuchi. The translation-rotation paradox in the nucleation theory. Journal of Statistical Physics, 1(2):351–375, ???? 1969. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01007485.

Knapp:1969:GFA

[KtH69] R. H. Knapp, Jr. and D. ter Haar. A Green function approach to the paramagnetic phase of a Heisenberg ferromagnet. Journal of Statistical Physics, 1(1):149–161, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01007247.

Mazo:1969:TBMa

[Maz69a] Robert M. Mazo. On the theory of Brownian motion. I. Interaction between Brownian particles. Journal of Statistical Physics, 1(1):89– 99, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.1007/
BF01007243.

Mazo:1969:TBMb

[Maz69b] Robert M. Mazo. On the theory of Brownian motion. II. Nonuniform systems. Journal of Statistical Physics, 1(1):101–106, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/BF01007244.

Mazo:1969:TBMc

[Maz69c] Robert M. Mazo. On the theory of Brownian motion. III. Two-body distribution function. Journal of Statistical Physics, 1(4):559–562, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.1007/
BF01024130.

Meecham:1969:SRN

[Mee69] W. C. Meecham. Stochastic representation of nearly-Gaussian, nonlinear processes. Journal of Statistical Physics, 1(1):25–40, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/BF01007239.

Mortensen:1969:MPM

[Mor69] Richard E. Mortensen. Mathematical problems of modeling stochastic nonlinear dynamic systems. Journal of Statistical Physics, 1(2):271– 296, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.1007/
BF01007481.

Narasimha:1969:VSC

[NDR69] R. Narasimha, S. M. Deshpande, and P. V. Subba Raju. Velocityspace contours of collision integrals. Journal of Statistical Physics, 1 (4):585–593, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01024133.

Nilsen:1969:NQD

[NH69] T. S. Nilsen and P. C. Hemmer. Note on the quantum displacement of the critical point. Journal of Statistical Physics, 1(1):175–178, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.1007/
BF01007249.

Orban:1969:EID

[OB69] J. Orban and A. Bellemans. Entropy and irreversibility in a dilute gas of hard disks. Journal of Statistical Physics, 1(3):467–474, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/BF01106581.

Posner:1969:DET

[PR69] Edward C. Posner and Eugene R. Rodemich. Differential entropy and tiling. Journal of Statistical Physics, 1(1):57–69, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:

//link.springer.com/article/10.1007/BF01007241.

Reiss:1969:I

[Rei69a] H. Reiss. Introduction. Journal of Statistical Physics, 1(1):1–2, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.1007/
BF01007237.

Reiss:1969:Pa

[Rei69b] H. Reiss. Prospectus. Journal of Statistical Physics, 1(3):377–381, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.1007/
BF01106577.

Reiss:1969:Pb

[Rei69c] H. Reiss. Prospectus. Journal of Statistical Physics, 1(4):601–608, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.1007/
BF01024136.

Reiss:1969:TLT

[Rei69d] H. Reiss. Thermodynamic-like transformations in information theory. Journal of Statistical Physics, 1(1):107–131, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:

//link.springer.com/article/10.1007/BF01007245.

Richardson:1969:PRS

[Ric69] John M. Richardson. Pattern recognition and statistical mechanics. Journal of Statistical Physics, 1(1):71–88, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:

//link.springer.com/article/10.1007/BF01007242.

Schilder:1969:NAW

[Sch69] Michael Schilder. New applications of Wiener integrals to engineering and physics. Journal of Statistical Physics, 1(3):475–516, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/BF01106582.

Seifer:1969:ISA

[Sei69] Arnold D. Seifer. On the incoherent scattering of an acoustic or electromagnetic wave. Journal of Statistical Physics, 1(4):571–584, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.1007/
BF01024132.

Stillinger:1969:LPG

[SS69] Frank H. Stillinger, Jr. and Zevi W. Salsburg. Limiting polytope geometry for rigid rods, disks, and spheres. Journal of Statistical Physics, 1(1):179–225, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01007250.

Simon:1969:CSU

[SZ69] Eliot M. Simon and Bruno H. Zimm. A computer simulation of the unwinding of a DNA-like helix. Journal of Statistical Physics, 1(1):41– 55, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.1007/
BF01007240.

Truesdell:1969:PUL

[Tru69] C. Truesdell. A precise upper limit for the correctness of the Navier– Stokes theory with respect to the kinetic theory. Journal of Statistical Physics, 1(2):313–318, ???? 1969. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.com/

article/10.1007/BF01007483.